OpenStudy (anonymous):

A fraction reduces to 36. If its denominator is 6x^5, what is its numerator?

OpenStudy (cwrw238):

multiply the denominator by 36

OpenStudy (anonymous):

It's 6^3x^5 right?

OpenStudy (cwrw238):

not quite - its 36 times 6x^5

OpenStudy (cwrw238):

- oh - you are also correct because 6*36 = 6^3 = 216

OpenStudy (cwrw238):

answer is therefore 216 x^5

OpenStudy (anonymous):

Thanks

OpenStudy (cwrw238):

yw